# ASBench: Image Anomalies Synthesis Benchmark for Anomaly Detection

### v1

@@ 4.4.3 Implications and Future Challenges
The limited predictive value for detection tasks of current evaluation requires the development of specialized assessment criteria for synthetic abnormal samples. An enhanced evaluation paradigm should concurrently operate at pixel-level and image-level dimensions, rather than rely solely on holistic image quality metrics. Such a dual-scale framework would enable more reliable quality assessment of synthetic anomalies while permitting perdiction when utilizing data for detector training, thereby conserving human and material resources typically expended in repetitive experimental iterations of detection tasks.
This challenge parallels difficulties encountered in evaluation of natural image generation. Some methods assess generation quality through combined global-local analyses. For example, portrait synthesis validation examines both overall reasonability and localized correctness. Drawing inspiration from these approaches, synthetic anomaly evaluation should incorporate Global consistency and Local fidelity. The establishment of such multidimensional evaluation standards represents a critical research direction for improving synthetic anomaly generation and application efficacy.
@@ 4.5.1 Comparative Experimental Framework
This section is based on four synthesis algorithms—Fractal, DRAEM, AnomalyDiffusion, and MemSeg—to construct and evaluate eleven hybrid combinations through training and testing, aiming to assess their synergistic effects in anomaly synthesis tasks. The experiments use the weighted average of performance metrics across the dataset as a reference. The results of the hybrid methods are compared with those of each individual method and theoretical combined means. Bar charts are plotted for different detection pipelines, as shown in Fig. 12, to analyze the performance differences among various combinations.
Figure 12: Comparison of anomaly detection performance (image-level AUROC) using mixed synthesis augmentation methods versus individual approaches. Blue part: Actual performance of hybrid synthesis methods. Yellow part: Theoretical performance computed from individual methods’ results. Red line: Maximum AUROC achieved by any single synthesis method.
4.5.2 Analysis of Results
From Fig. 12, it can be observed that most hybrid methods outperform the best results achieved by individual methods. A few hybrid methods show slightly lower performance than the optimal single method, but still surpass the weaker individual methods. Additionally, by comparing the theoretical and actual performance of hybrid methods, we find that the actual performance of hybrid methods generally exceeds the theoretical mean, indicating that the combination of multiple methods indeed produces synergistic effects. And the comparison of visualization is shown in Fig. 13.
@@ 4.5.2 Analysis of Results
Notably, mixed methods incorporating AnomalyDiffusion show significant improvement in the appixel metric compared to others. AnomalyDiffusion generates abnormal samples with high quality and diversity at pixel-level, better simulating real-world anomaly data. This allows the model to learn more refined anomaly features during training, while other methods primarily focus on image-level.
@@ 4.5.3 Implications and Future Challenges
Although the diversity, complementarity, and robustness of different methods are advantages of hybrid approaches, selecting optimal combination of methods and adjusting their weights remains a complex optimization problem. The significant improvement in pixel-level performance when AnomalyDiffusion is incorporated highlights necessity of including high-quality pixel-level samples in anomaly synthesis. In other words, it is worth exploring anomaly synthesis methods that operate at both image-level and pixel-level.
@@ 5 Conclusion
We introduce ASBench in this paper, the first comprehensive benchmark for image anomaly synthesis, featuring 12 anomaly synthesis methods, 4 anomaly detection pipelines, and 5 industrial datasets across 4 key evaluation dimensions. Through extensive experiments, we have gained crucial insights into the performance of anomaly synthesis, including the discovery that no single method is universally optimal; the non-linear influence of synthetic data ratios, where higher proportions do not guarantee better performance; the weak correlation between conventional image quality metrics and downstream detection efficacy; and the notable performance boost from hybridizing complementary synthesis methods. On top of these findings, we present several intriguing future lines for image anomaly synthesis.

### v3

@@ 4.4.3 Implications and Future Challenges
The limited predictive value for detection tasks of current evaluation requires the development of specialized assessment criteria for synthetic abnormal samples. An enhanced evaluation paradigm should concurrently operate at pixel-level and image-level dimensions, rather than rely solely on holistic image quality metrics. Such a dual-scale framework would enable more reliable quality assessment of synthetic anomalies while permitting prediction when utilizing data for detector training, thereby conserving human and material resources typically expended in repetitive experimental iterations of detection tasks.
This challenge parallels difficulties encountered in evaluation of natural image generation. Some methods assess generation quality through combined global-local analyzes. For example, portrait synthesis validation examines both overall reasonability and localized correctness. Drawing inspiration from these approaches, synthetic anomaly evaluation should incorporate Global consistency and Local fidelity. The establishment of such multidimensional evaluation standards represents a critical research direction for improving synthetic anomaly generation and application efficacy.
@@ 4.5.2 Comparative Experimental Framework
This section is based on four synthesis algorithms—Fractal, DRAEM, AnomalyDiffusion, and MemSeg—to construct and evaluate eleven hybrid combinations through training and testing, aiming to assess their synergistic effects in anomaly synthesis tasks. The experiments use the weighted average of performance metrics across the dataset as a reference. The results of the hybrid methods are compared with those of each individual method and theoretical combined means. The comparison of visualization is shown in Fig. 12. And bar charts are plotted for different detection pipelines, as shown in Fig. 13, to analyze the performance differences among various combinations.
Figure 13: Comparison of performance (image-level AUROC) using mixed synthesis augmentation methods versus individual approaches. Blue part: Actual performance of hybrid synthesis methods. Yellow part: Theoretical performance computed from individual methods’ results. Red line: Maximum AUROC achieved by any single synthesis method.
From Fig. 13, it can be observed that most hybrid methods outperform the best results achieved by individual methods. A few hybrid methods show slightly lower performance than the optimal single method, but still surpass the weaker individual methods. Additionally, by comparing the theoretical and actual performance of hybrid methods, we find that the actual performance of hybrid methods generally exceeds the theoretical mean, indicating that the combination of multiple methods indeed produces synergistic effects.
@@
Notably, mixed methods incorporating AnomalyDiffusion show significant improvement in the pixel-level metric compared to others. AnomalyDiffusion generates abnormal samples with high quality and diversity at pixel-level, better simulating real-world anomaly data. This allows the model to learn more refined anomaly features during training, while other methods primarily focus on image-level.
@@ 4.5.3 Implications and Future Challenges
Although the diversity, complementarity, and robustness of different methods are advantages of hybrid approaches, selecting optimal combination of methods and adjusting their weights remains a complex optimization problem. The significant improvement in pixel-level performance when AnomalyDiffusion is incorporated highlights necessity of including high-quality pixel-level samples. In other words, it is worth exploring anomaly synthesis methods that operate at both image and pixel level.
@@ 5 Conclusion
We introduce ASBench in this paper, the first comprehensive benchmark for image anomaly synthesis, featuring 12 anomaly synthesis methods, 4 anomaly detection pipelines, and 5 industrial datasets across 4 key evaluation dimensions. Through extensive experiments, we have gained crucial insights into the performance of anomaly synthesis, including the discovery that no single method is universally optimal; the non-linear influence of synthetic data ratios, where higher proportions do not guarantee better performance; the weak correlation between conventional image quality metrics and downstream detection efficacy; and the notable performance boost from hybridizing complementary synthesis methods. On top of these findings, we present several intriguing future lines for image anomaly synthesis. For example, future work could focus on (1) designing anomaly synthesis methods with greater generalizability to balance diversity and realism; (2) developing adaptive synthesis strategies to optimize the trade-off between the quantity and quality of synthetic samples; (3) constructing novel task-oriented evaluation metrics that can reliably predict detection performance; and (4) systematically optimizing the combination and weighting mechanisms when hybridizing multiple synthesis approaches.